# Direct measurement of atomic number density using Single Pass Absorption Spectroscopy (SPAS)

### v1

@@ Acknowledgments
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@
The authors declare no conflict of interest.
Appendix A Measurement of transmitted power after commercial vapor cell
The transmitted laser intensity through the rubidium vapor cell was measured as a function of the incident intensity using a calibrated power meter. Figure 6 compares the experimental data with theoretical predictions obtained from the Beer–Lambert law. In the theoretical model, attenuation due to the glass cell walls was independently measured in the off-resonant region and incorporated into the transmission calculation.
Figure 6: Comparison of experimental and theoretical transmitted laser intensity through the rubidium vapor cell of length 100 mm at a temperature of $20^{\circ}$C. The blue stars represent the measured transmitted intensity obtained using a calibrated power meter, while the red circles with connecting lines indicate the theoretical values calculated from our model using the Beer–Lambert law with the inclusion of off-resonant glass cell wall losses.
For a cell temperature of $20^{\circ}$C, the measured transmission follows the theoretical expectation across the full range of incident intensities investigated. The linear correlation observed between incident ($I_{0}$) and transmitted ($I(z)$) intensities confirms that absorption remains in the weak-probe regime, where the Beer–Lambert formalism accurately describes the system. The excellent agreement between experiment and theory indicates that both the vapor absorption and the cell wall losses have been accounted for in the model, thereby validating the use of absolute absorption spectroscopy for extracting quantitative number density information under these conditions.
Appendix B Measurement of beam diameter of the laser
The Rabi frequency plays a very important role in estimating the absolute absorption profile of alkali atoms. To accurately calculate the Rabi frequency, we need to know the beam diameter, as the Rabi frequency inversely depends on the beam’s cross-sectional area. Thus, a precise measurement of the laser beam diameter is very crucial. As the spatial profile of our laser beam is elliptical, the best way to measure the beam diameter ($1/e^{2}$ diameter) is using the centroid method. To accurately determine the beam dimensions, we captured its image using a CCD camera (with a pixel size of 2.8 $\mu m$) at a position just before the vapor cell, as illustrated in Figure 7.
The recorded beam profile was analyzed according to the ISO 11146 standard [34], which prescribes the use of second-order statistical moments to determine the beam width and orientation. Specifically, first-order moments were used to identify the centroid of the beam (center of mass), while second-order moments provided information about the spatial variance and the azimuthal tilt of the beam [35]. The beam diameters corresponding to the $1/e^{2}$ intensity level were extracted along the principal axes of the elliptical profile, which were rotated with respect to the CCD pixel grid. These diameters, initially calculated in pixel units, were converted to physical dimensions using the known pixel size. Based on this analysis, the beam diameter along the major axis was measured to be $2.22\pm 0.04$ mm, while the minor axis diameter was found to be $1.52\pm 0.03$ mm with the rotation angle of the principal axis 4.15.
Appendix C Absorption spectra at high power
@@
We tried an iterative approach that accounts for the intensity-dependent susceptibility along the propagation path to accurately model absorption at these higher intensities. Rather than treating the vapor cell as a uniform medium, we divided the 100 mm cell into $n$ equal slices along the laser propagation axis, where each slice can be considered to have constant incident laser intensity along its length. The number of slices $n$ can be adjusted based on computational resources and desired accuracy, with more slices providing higher precision at the cost of increased computation time.
@@
where $\chi(I(z),\Delta,T)$ is the electric susceptibility, which mainly depends on the temperature (T) of the vapor cell, the detuning ($\Delta$) of the laser, and the local beam intensity ($I(z)$). This intensity dependence captures the saturation effects that become prominent at higher powers, allowing accurate modeling of the non-linear absorption process beyond the weak probe limit.
@@
The strong agreement across different probe powers, from the linear weak-probe regime to the onset of saturation, confirms the reliability and versatility of our model. This validation supports its use for extracting accurate number density of alkali vapors and transition parameters beyond the weak probe regime.
Appendix D Characterization of number density for MEMS vapor cell
The number density of rubidium in the MEMS cell was extracted from absorption measurements over a range of temperatures and compared to the empirical relation from Equation 12. As shown in Figure 9, the experimental data follow the theoretical temperature dependence, validating the applicability of vapor pressure models even in microscale geometries. The insets highlight specific regions at $25^{\circ}$C, $40^{\circ}$C, and $60^{\circ}$C, showing consistency between the measured and predicted values within the experimental uncertainty. These results confirm that our model can accurately measure the number density of a compact MEMS vapor cell. The successful application of the model to MEMS cells has important implications for the development of chip-scale atomic devices, where precise knowledge of the atomic density is crucial for optimizing sensor performance. The ability to accurately determine vapor density in these miniaturized systems enables better calibration and characterization of MEMS-based atomic sensors, quantum sensors, and precision timing devices.
Figure 9: Rubidium vapor number density as a function of temperature, showing experimental MEMS cell data compared to the empirical model predictions. Insets (a), (b), and (c) provide detailed views at 25°C, 40°C, and 60°C, respectively.

### v2

@@ Acknowledgments
The authors declare no conflicts of interest.
Appendix A Characterization of number density for 100 mm Rb vapor cell
We also determine the rubidium vapor number density in a 100 mm long vapor cell using absorption spectroscopy and quantitative spectral fitting. This analysis establishes a reliable, self-consistent characterization of the rubidium vapor number density in the 100 mm cell. The experimental schematic for the single-pass absorption spectroscopy is shown in Figure 7.
Figure 7: Schematic of the experimental setup for absolute absorption spectroscopy and frequency calibration for a 100 mm long vapor cell.
Figure 8 shows the measured absorption spectra as a function of laser detuning at three cell temperatures: $293.45~\pm~0.3$ K, $312.65~\pm~0.8$ K, and $339.65~\pm~3.3$ K. We fit the experimental spectra using a model that includes Doppler broadening, natural linewidth, and temperature-dependent atomic number density. The fitted spectra accurately reproduce the peak absorption, relative transition strengths, and off-resonant wings at all temperatures. The residuals (experiment minus fit), shown in the lower panels of Figure 8, remain within $\pm 0.25$ in absolute transmission.
Figure 8: Experimental and fitted absorption spectra as a function of detuning for three different temperatures: 293.45 K (red solid line for fit, blue dots for experiment), 312.65 K (black dashed line for fit, orange dots for experiment), and 339.65 K (blue dotted line for fit, green dots for experiment). The lower panels show the corresponding residuals (experimental minus fitted). The fitted spectra exhibit excellent agreement with experimental data across all temperatures, with residuals remaining within $\pm~0.25$ in absolute transmission.
From these fits, we extract the rubidium vapor number density as a function of temperature. Figure 10 shows the extracted number densities obtained independently from absorption measurements at 780 nm and 420 nm, together with the empirical rubidium vapor pressure model. The two experimental datasets show strong mutual agreement and closely follow the model over the full temperature range. The vertical dashed line at 312.45 K marks the solid–liquid phase transition of rubidium, where the temperature dependence of the number density changes slope, consistent with thermodynamic expectations. The insets highlight the agreement between experiment and model at selected temperatures.
Figure 9: Comparison of simulated and experimental transmission spectra of the 100 mm vapor cell at 310.65 K and 331.75 K. Solid lines represent the simulated spectra, and markers show the corresponding experimental data. The lower panels display the absolute residuals (experiment – simulation) for each temperature.
We further validate the extracted number densities by directly comparing simulated and experimental transmission spectra. Figure 9 presents this comparison at 310.65 K and 331.75 K using the corresponding fitted densities. The simulated spectra reproduce both the magnitude and spectral shape of the measured transmission. The residuals remain small across the full frequency scan, confirming that the extracted number densities provide a quantitatively accurate description of the rubidium vapor in the 100 mm cell.
Figure 10: Rubidium vapor number density as a function of temperature for a 100 mm vapor cell. Experimental number densities extracted from absorption measurements at 780 nm (red circles) and 420 nm (blue diamonds) are compared with the empirical vapor pressure model (dashed line). The vertical dashed line marks the phase transition point at 312.45 K. Insets (a), (b), and (c) show enlarged views of the data around 293 K, 313 K, and 341 K, respectively.
This analysis establishes a reliable, self-consistent characterization of the rubidium vapor number density in the 100 mm cell, providing a robust basis for subsequent precision spectroscopy and temperature-dependent studies.
To further validate the generality of the model, absorption spectroscopy was also performed on isotopically enriched rubidium vapor cells: an $86.5\%$ enriched 85Rb cell and a $93.6\%$ enriched 87Rb cell, each with a length of $100$ mm. Spectra were recorded at the same three cell temperatures ($293.45~\pm~0.3$ K, $312.65~\pm~0.8$ K, and $339.65~\pm~3.3$ K). For both isotopic compositions, the experimental spectra were well reproduced by the theoretical model, yielding $R^{2}>0.99$ in all cases. This consistency highlights the model’s versatility and reliability across different isotopic mixtures.
Although the NTC temperature sensors used in our experiment setup have an intrinsic accuracy of $\pm~0.3$ K, the observed error bars in the number density of the 100 mm long vapor cell at higher temperatures are larger than at lower temperatures. At higher temperatures, we observe that rubidium atoms in the vapor cell tend to deposit on the inner walls through which the laser beam passes. This introduces abrupt non-linearity in the absorption signal. To mitigate this issue, we introduced a slight temperature gradient along the cell, so that the flat glass window through which the laser beam passes remains slightly hotter than the center cylindrical region. Although this configuration helps us suppress the condensation on the flat window, it marginally increases the uncertainty in the temperature measurement.
Appendix B Absorption spectra at high power
@@
We used an iterative approach that accounts for intensity-dependent susceptibility along the propagation path to accurately model absorption at higher intensities. Rather than treating the vapor cell as a uniform medium, we divided the 100 mm cell into $n$ equal slices along the laser propagation axis, where each slice can be considered to have constant incident laser intensity along its length. The number of slices $n$ can be adjusted based on computational resources and desired accuracy, with more slices providing higher precision at the cost of increased computation time.
@@
where $\chi(I(z),\Delta,T)$ is the electric susceptibility, which mainly depends on the temperature (T) of the vapor cell, the detuning ($\Delta$) of the laser, and the local beam intensity ($I(z)$). This intensity dependence captures the saturation effects that become prominent at higher powers, allowing accurate modeling of the nonlinear absorption process beyond the weak-probe limit.
@@
The strong agreement across different probe powers, from the linear weak-probe regime to the onset of saturation, confirms the reliability and versatility of our model. This validation supports its use for extracting accurate number density of alkali vapors and transition parameters beyond the weak-probe regime.
Appendix C Measurement of laser beam diameter
[…]
The recorded beam profile was analyzed according to the ISO 11146 standard [48], which prescribes the use of second-order statistical moments to determine the beam width and orientation. Specifically, first-order moments were used to identify the centroid of the beam (center of mass), while second-order moments provided information about the spatial variance and the azimuthal tilt of the beam [49]. The beam diameters corresponding to the $1/e^{2}$ intensity level were extracted along the principal axes of the elliptical profile, which were rotated with respect to the CCD pixel grid. These diameters, initially calculated in pixel units, were converted to physical dimensions using the known pixel size. Based on this analysis, the beam diameter along the major axis was measured to be $2.22~\pm~0.04$ mm, while the minor axis diameter was found to be $1.52~\pm~0.03$ mm with the rotation angle of the principal axis 4.15.
Appendix D Measurement of transmitted power after commercial vapor cell
The transmitted laser intensity through the rubidium vapor cell was measured as a function of the incident intensity using a calibrated power meter. Figure 13 compares the experimental data with theoretical predictions obtained from the Beer–Lambert law. In the theoretical model, attenuation due to the glass cell walls was independently measured in the off-resonant region and incorporated into the transmission calculation. For a cell temperature of 293.15 K, the measured transmission follows the theoretical expectation across the full range of incident intensities investigated. The linear correlation observed between incident ($I_{0}$) and transmitted ($I(z)$) intensities confirms that absorption remains in the weak-probe regime, where the Beer–Lambert formalism accurately describes the system. The excellent agreement between experiment and theory indicates that both the vapor absorption and the cell wall losses have been accounted for in the model, thereby validating the use of absolute absorption spectroscopy for extracting quantitative number density information under these conditions.
Figure 13: Comparison of experimental and theoretical transmitted laser intensity through the rubidium vapor cell of length 100 mm at a temperature of 293.15 K. The blue stars represent the measured transmitted intensities obtained with a calibrated power meter, while the red circles with connecting lines indicate the theoretical values calculated from our model using the Beer–Lambert law, including off-resonant glass-cell wall losses.